OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. S = 5r²t, for t

OpenStudy (wwe123):

t = S/5r^2

OpenStudy (anonymous):

how did you figure tht out i need help on figuring it out not just the answer

OpenStudy (dape):

Well you want to isolate t, that means you want everything except t to be on one side of the equals-sign. The method for that is to do the reverse operations wherever t is not alone, so in this t is multiplied by \(5r^2\) so we divide both sides by \(5r^2\), and the result is \[ \frac{S}{5r^2}=\frac{5r^2t}{5r^2} \] On the right side we can cancel \(5r^2\), so the answer is \[ t=\frac{S}{5r^2} \]